### v2

# On the $\Gamma$-limit of weighted fractional energies

Abstract.
Given $p\in[1,\infty)$ and a bounded open set $\Omega\subset\mathbb{R}^{d}$ with Lipschitz boundary, we study the $\Gamma$-convergence of the weighted fractional seminorm
$u]_{s,p,f}^{p}=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|\tilde{u}(x)-\tilde{u}(y)|^{p}}{\|x-y\|^{d+sp}}\,f(x)\,f(y)\,\mathrm{d}x\,\mathrm{d}y,$
as $s\to 1^{-}$ for $u\in L^{p}(\Omega)$, where $\tilde{u}=u$ on $\Omega$ and $\tilde{u}=0$ on $\mathbb{R}^{d}\setminus\Omega$. Assuming that $(f_{s})_{s\in(0,1)}\subset L^{\infty}(\mathbb{R}^{d};[0,\infty))$ and $f\in\mathrm{Lip}_{b}(\mathbb{R}^{d};(0,\infty))$ are such that $f_{s}\to f$ in $L^{\infty}(\mathbb{R}^{d})$ as $s\to 1^{-}$, we show that $(1-s)[u]_{s,p,f_{s}}^{p}$ $\Gamma$-converges to the Dirichlet $p$-energy weighted by $f^{2}$. In the case $p=2$, we also prove the convergence of the corresponding gradient flows.
Key words and phrases:
$\Gamma$-convergence, fractional gradient flows, Gagliardo seminorms, parabolic flows, weighted spaces
2020 Mathematics Subject Classification:
Primary 49J45. Secondary 35R11, 35B40, 45E10, 26A33
Acknowledgements. The second-named and third-named authors are members of the Istituto Nazionale di Alta Matematica (INdAM), Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA). The first-named author is supported by the Academy of Finland (grant agreement No. 314227). The second-named author has received funding from INdAM under the INdAM–GNAMPA Project 2025 Disuguaglianze funzionali di tipo geometrico e spettrale (grant agreement No. CUP_E5324001950001). The third-named author has received funding from INdAM under the INdAM–GNAMPA Project 2025 Metodi variazionali per problemi dipendenti da operatori frazionari isotropi e anisotropi (grant agreement No. CUP_E5324001950001) and from the European Union – NextGenerationEU and the University of Padua under the 2023 STARS@UNIPD Starting Grant Project New Directions in Fractional Calculus – NewFrac (grant agreement No. CUP_C95F21009990001).
This is a pre-print of an article published in Proc. Roy. Soc. Edinburgh Sect. A. The final authenticated version is available online at: https://doi.org/10.1017/prm.2025.10100
1. Introduction
1.1. Framework
We let $d\in\mathbb{N}$, $s\in(0,1)$, and $p\in[1,\infty)$. Given a nonnegative weight $f\in L^{\infty}(\mathbb{R}^{d};[0,\infty))$, our aim is to study the $\Gamma$-convergence as $s\to 1^{-}$ of the non-homogeneous (or weighted by $f$) $s$-fractional $p$-seminorm
$u]_{s,p,f}^{p}=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{\|x-y\|^{d+sp}}\,f(x)\,f(y)\,\mathrm{d}x\,\mathrm{d}y,$ (1.1)
for $u\in L^{p}(\mathbb{R}^{d})$.
The convergence as $s\to 1^{-}$ of (1.1) in the case $f\equiv 1$—for which we use the shorthand $\cdot\,]_{s,p}^{p}=[\,\cdot\,]_{s,p,1}^{p}$—has been deeply studied in recent years, both in the pointwise and in the $\Gamma$-sense. Since the literature is very vast, here we limit ourselves to a non-comprehensive list of results which are closer to the spirit of the present work.
The pointwise limit of the seminorm $\cdot\,]^{p}_{s,p}$ as $s\to 1^{-}$ is a notable instance of the celebrated Bourgain–Brezis–Mironescu (BBM, for short) formula [BBM01, Dav02], yielding that $(1-s)[\,\cdot\,]_{s,p}$ converges to the Dirichlet $p$-energy up to a multiplicative constant. After the seminal contributions [BBM01, Dav02], the BBM formula has been extensively studied in several directions, see [Pon04a, Pon04b, DDG24] for more general results and [LS11, LS14] for extensions to arbitrary domains. We also refer to [Fan25, Lud14-S, Lud14-p] for anisotropic fractional energies and to [DDP24, GS25] for sharp conditions for the validity of the BBM formula.
The $\Gamma$-convergence of $(1-s)[\,\cdot\,]_{s,p}^{p}$ to the Dirichlet $p$-energy as $s\to 1^{-}$ has been established in [BPS16] for every $p\in(1,\infty)$, in [CdLKNP23] only for $p=2$, and in [GS25] for every $p\in[1,\infty)$. We also refer to [BBM01, Pon04a] for similar results on bounded open sets.
The geometric case $p=1$ deserves special mention, due to the link with the (relative) fractional perimeter, see [AdPM01, P20, BP19, KlM24, GS25, Lom19] for closely related results in this direction. We also refer to [CN20, dLKP22, KPT2024] for higher-order convergence results.
Beyond the case $f\equiv 1$, the asymptotic behavior of (1.1) and of similarly-defined energies has been studied for some particular weights, see [CCLP23, DL21] for the Gaussian framework and [K24] for weights depending on negative powers of the distance from the boundary.
The aim of the present paper is to investigate the asymptotic behavior of the weighted seminorm (1.1) as $s\to 1^{-}$ as the weight $f$ is also allowed to vary with respect to the parameter $s$. This is motivated by the recent interest in the extension of BBM-type formulas beyond the isotropic setting in order to address possible applications to non-isotropic frameworks [DDP24]. Besides, our $\Gamma$-convergence result can be interpreted as a suitable extension to the weighted setting of the ones obtained in [CdLKNP23, BPS16].
Our main result, Theorem˜1.1 below, deals with the $\Gamma$-convergence of the energy (1.1) with respect to a uniformly converging family of weights $(f_{n})_{n\in\mathbb{N}}$ in $L^{\infty}(\mathbb{R}^{d};[0,\infty))$, whose limit $f$ is in $\mathrm{Lip}_{b}(\mathbb{R}^{d};(0,\infty))$. Precisely, we prove that the $\Gamma$-limit is given by
$u\mapsto\begin{cases}\displaystyle K_{d,p}\|\nabla u\|_{p,f^{2}}^{p}=K_{d,p}\int_{\Omega}f^{2}\,\|\nabla u\|^{p}\,\mathrm{d}x,&\text{for}\ p\in(1,\infty),\\[12.91663pt] \displaystyle K_{d,1}\|Du\|_{1,f^{2}}=K_{d,1}\int_{\Omega}f^{2}\,\mathrm{d}|Du|,&\text{for}\ p=1,\end{cases}$
where, for every $p\in[1,\infty)$ (and here $\Gamma$ being Euler’s Gamma function),
$K_{d,p}=\frac{1}{p}\int_{\partial B_{1}}|x\cdot\mathrm{e}_{d}|^{p}\,\mathrm{d}\mathcal{H}^{d-1}(x)=\frac{2\pi^{\frac{d-1}{2}}}{p}\frac{\Gamma\left(\frac{p+1}{2}\right)}{\Gamma\left(\frac{N+p}{2}\right)},$ (1.2)
see [BSZ26]*Lem. 2.1. Here and below, given a measurable function $u\colon\Omega\to\mathbb{R}$ on an open set $\Omega\subset\mathbb{R}^{d}$, we define $\tilde{u}\colon\mathbb{R}^{d}\to\mathbb{R}$ be such that $\tilde{u}=u$ on $\Omega$ and $\tilde{u}=0$ on $\mathbb{R}^{d}\setminus\Omega$.
Theorem 1.1 ($\Gamma$-convergence with weights).
Let $p\in[1,\infty)$, $(f_{n})_{n\in\mathbb{N}}\subset L^{\infty}(\mathbb{R}^{d};[0,\infty))$ and $f\in\mathrm{Lip}_{b}(\mathbb{R}^{d};(0,\infty))$ be such that $f_{n}\to f$ in $L^{\infty}(\mathbb{R}^{d})$, $\Omega\subset\mathbb{R}^{d}$ be a bounded open set with Lipschitz boundary and $(s_{n})_{n\in\mathbb{N}}\subset(0,1)$ be such that $s_{n}\to 1^{-}$.
  1. (i)
    (Compactness) If $(u^{n})_{n\in\mathbb{N}}\subset L^{p}(\Omega)$ is such that
    $\sup_{n\in\mathbb{N}}\left((1-s_{n})[\tilde{u}^{n}]^{p}_{s_{n},p,f_{n}}+\|u^{n}\|_{L^{p}(\Omega)}^{p}\right)<\infty,$ (1.3)
    then, up to a subsequence, $u^{n}\to u$ in $L^{p}(\Omega)$ for some $u\in W^{1,p}_{0}(\Omega)$ if $p>1$ or $u\in BV(\Omega)$ if $p=1$.
  2. (ii)
    ($\Gamma$-liminf inequality) If $(u^{n})_{n\in\mathbb{N}}\subset L^{p}(\Omega)$ is such that $u^{n}\to u$ in $L^{p}(\Omega)$ for some $u\in W^{1,p}_{0}(\Omega)$ if $p>1$, or $u\in BV(\Omega)$ if $p=1$, then
    $\begin{split}&K_{d,p}\|\nabla u\|^{p}_{p,f^{2}}\leq\liminf_{n\to\infty}(1-s_{n})[\tilde{u}^{n}]^{p}_{s_{n},p,f_{n}}\quad\text{for $p>1$},\\ &K_{d,1}\|Du\|_{1,f^{2}}\leq\liminf_{n\to\infty}(1-s_{n})[\tilde{u}^{n}]_{s_{n},1,f_{n}}\quad\text{for $p=1$}.\end{split}$ (1.4)
  3. (iii)
    ($\Gamma$-limsup inequality) If $u\in W^{1,p}_{0}(\Omega)$ if $p>1$, or $u\in BV(\Omega)$ if $p=1$, then there exists $(u^{n})_{n\in\mathbb{N}}\subset L^{p}(\Omega)$ such that $u^{n}\to u$ in $L^{p}(\Omega)$ and
    $\begin{split}&K_{d,p}\|\nabla u\|^{p}_{p,f^{2}}=\lim_{n\to\infty}(1-s_{n})[\tilde{u}^{n}]^{p}_{s_{n},p,f_{n}}\quad\text{for $p>1$},\\ &K_{d,1}\|Du\|_{1,f^{2}}=\lim_{n\to\infty}(1-s_{n})[\tilde{u}^{n}]_{s_{n},1,f_{n}}\quad\text{for $p=1$}.\end{split}$ (1.5)
1.2. Convergence of flows
In the case $p=2$, the $\Gamma$-convergence result obtained in Theorem˜1.1 can be complemented with a stability result for the corresponding parabolic flows associated to the energies, see Theorem˜1.2 below. Here and below, given a weight $f\in L^{\infty}(\mathbb{R}^{d};[0,\infty))$, we define the weighted Laplacian of $u\in H^{1}_{0}(\Omega)$ as
$(-\mathfrak{D})^{f}u=-2\operatorname{div}(f^{2}\nabla u)$ (1.6)
in the distributional sense in duality with $C^{\infty}_{c}(\Omega)$ functions. Moreover, given $u\in L^{2}(\Omega)$ such that $u]_{s,2,f}<\infty$, we define the weighted fractional $s$-Laplacian of $u$ as
$(-\mathfrak{D})^{s,f}u(x)=4f(x)\,\mathrm{p.v.}\int_{\mathbb{R}^{d}}\frac{\tilde{u}(x)-\tilde{u}(y)}{\|x-y\|^{d+2s}}\,f(y)\,\mathrm{d}y$ (1.7)
again in the distributional sense in duality with $C^{\infty}_{c}(\Omega)$ functions. Note that, in the unweighted case $f\equiv 1$, up to a multiplicative constant, the operators (1.6) and (1.7) become the usual Laplacian and fractional $s$-Laplacian operators, respectively.
Theorem 1.2 (Stability of parabolic flows).
Let $(f_{n})_{n\in\mathbb{N}}$, $f$, $\Omega$, and $(s_{n})_{n\in\mathbb{N}}$ be as in Theorem˜1.1. If $(u^{n}_{0})_{n\in\mathbb{N}}\subset L^{2}(\Omega)$ is such that $u^{n}_{0}\to u^{\infty}_{0}$ in $L^{2}(\Omega)$ for some function $u^{\infty}_{0}\in L^{2}(\Omega)$, $\tilde{u}^{n}_{0}]_{s_{n},2,f_{n}}<\infty$ for every $n\in\mathbb{N}$, and
$\sup_{n\in\mathbb{N}}(1-s_{n})[\tilde{u}^{n}_{0}]_{s_{n},2,f_{n}}^{2}<\infty,$
then the following hold:
  1. (i)
    $u^{\infty}_{0}\in H^{1}_{0}(\Omega)$;
  2. (ii)
    for every $T>0$ and for every $n\in\mathbb{N}$, the problem
    $\begin{dcases}\dot{u}(t)=(1-s_{n})(-\mathfrak{D})^{s_{n},f_{n}}u(t),\quad\textrm{for a.e.\ }t\in(0,T),\\[4.30554pt] u(0)=u^{n}_{0},\end{dcases}$
    admits a unique solution $u_{n}\in H^{1}([0,T];L^{2}(\Omega))$ such that
[…]
  3. (iii)
    the problem
    $\begin{dcases}\dot{u}(t)=K_{d,2}(-\mathfrak{D})^{f}u(t),\quad\text{for a.e.}\ t\in[0,\infty),\\[4.30554pt] u(0)=u^{\infty}_{0},\end{dcases}$
    admits a unique solution $u_{\infty}\in H^{1}([0,T];H^{1}_{0}(\Omega))$;
  4. (iv)
    $(u_{n})_{n\in\mathbb{N}}$ weakly converges to $u_{\infty}$ in $H^{1}([0,T];L^{2}(\Omega))$.
Moreover, if
$\lim_{n\to\infty}(1-s_{n})[\tilde{u}^{n}_{0}]^{2}_{s_{n},2,f_{n}}=K_{d,2}\|\nabla u^{\infty}_{0}\|^{2}_{2,f^{2}},$
then $(u_{n})_{n\in\mathbb{N}}$ strongly converges to $u_{\infty}$ in $H^{1}([0,T];L^{2}(\Omega))$ and also
$u_{n}(t)\xrightarrow{L^{2}}u_{\infty}(t)\quad\text{and}\quad(1-s_{n})[\tilde{u}_{n}(t)]_{s_{n},2,f_{n}}\to K_{d,2}\|\nabla u_{\infty}(t)\|_{2,f^{2}}\quad\text{for every}\ t\in[0,T].$
1.3. Organization of the paper
The paper is organized as follows. The notation and some useful preliminary results are detailed in Section˜2. The proof of Theorem˜1.1 is given in Section˜3, while that of Theorem˜1.2 can be found in Section˜4.
2. Preliminaries
2.1. Notation
We briefly detail the main notation used throughout the paper.
The symbol $C(*,\cdots,*)$ indicates a generic positive constant that depends on $\cdots,*$ only and may change from line to line.
We let $d\in\mathbb{N}$ and work in the $d$-dimensional Euclidean space $\mathbb{R}^{d}$. We let $x\cdot y$ be the Euclidean inner product between $x,y\in\mathbb{R}^{d}$ and $\|x\|$ be the Euclidean norm of $x$.
We let $B_{r}(x)$ be the open ball in $\mathbb{R}^{d}$ of center $x\in\mathbb{R}^{d}$ and radius $r>0$, and we use the shorthand $B_{r}=B_{r}(0)$. Given an open set $A\subset\mathbb{R}^{d}$, we let $A^{c}=\mathbb{R}^{d}\setminus A$ be the complement of $A$, $\partial A$ be the topological boundary of $A$ and, for every $t>0$,
[…]
Throughout the paper, we let $\Omega\subset\mathbb{R}^{d}$ be a bounded open set with Lipschitz boundary.
We let $\mathcal{L}^{d}$ be the $d$-dimensional Lebesgue measure and $\mathcal{H}^{\alpha}$ be the $\alpha$-dimensional Hausdorff measure for every $\alpha\in[0,d]$. We set $\omega_{d}=\mathcal{L}^{d}(B_{1})$, so that $\mathcal{H}^{d-1}(\partial B_{1})=d\,\omega_{d}$. Throughout the paper, all functions and sets are tacitly assumed to be $\mathcal{L}^{d}$-measurable.
Let $p\in[1,\infty)$ and $f\in L^{\infty}(\mathbb{R}^{d};[0,\infty))$. Given $m\in\mathbb{N}$ and $v\colon\Omega\to\mathbb{R}^{m}$, we let
$\|v\|_{p,f}=\left(\int_{\Omega}\|v(x)\|^{p}\,f(x)\,\mathrm{d}x\right)^{\frac{1}{p}}\in[0,\infty],$ (2.2)
and we use the shorthand $\|v\|_{p}=\|v\|_{p,1}$. We thus let
[…]
When $m=1$, we simply write $L^{p}_{f}(\Omega)$. We point out that if additionally $f$ takes values in $(0,\infty)$, under our standing assumptions on $f$ and $\Omega$, the spaces $L^{p}(\Omega)$ and $L^{p}_{f}(\Omega)$ are equivalent, with
$(\operatorname{ess\,inf}_{\Omega}f)\,\|v\|_{p}^{p}\leq\|v\|^{p}_{p,f}\leq\|f\|_{\infty}\|v\|^{p}_{p}.$ (2.3)
Given $u\colon\Omega\to\mathbb{R}$, we define $\tilde{u}\colon\mathbb{R}^{d}\to\mathbb{R}$ by letting $\tilde{u}=u$ in $\Omega$ and $\tilde{u}=0$ in $\mathbb{R}^{d}\setminus\Omega$. Thus, given $s\in(0,1)$, we define
$u]_{s,p,f}=\left(\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|\tilde{u}(x)-\tilde{u}(y)|^{p}}{\|x-y\|^{d+sp}}\,f(x)\,f(y)\,\mathrm{d}x\,\mathrm{d}y\right)^{\frac{1}{p}}$ (2.4)
for every $u\colon\Omega\to\mathbb{R}$ and we use the shorthand $u]_{s,p}=[u]_{s,p,1}$.
Finally, we let $W^{1,p}_{0}(\Omega)$ for $p>1$, be the closure of $C^{\infty}_{c}(\Omega)$ functions with respect to the Sobolev $p$-norm $u\mapsto\|u\|_{p}^{p}+\int_{\mathbb{R}^{d}}\|\nabla u\|_{p}^{p}\,\mathrm{d}x$, while $BV(\Omega)$ the weak closure of $C^{\infty}_{c}(\Omega)$ with respect to the Sobolev $1$-norm. We also set
$\|Du\|_{1,f}=\int_{\mathbb{R}^{d}}f\,\mathrm{d}|Du|,$ (2.5)
whenever $u\in BV(\Omega)$ and $f\in L^{\infty}(\mathbb{R}^{d};(0,\infty))$.
2.2. Compactness and characterization
We recall the following well-known compactness result, see [Bre11book]*Thm. 4.26 for example. Here and below, we let $\tau_{h}w(\,\cdot\,)=w(\,\cdot\,+h)$ for every $h\in\mathbb{R}^{d}$ and $w\in L^{p}(\mathbb{R}^{d})$.
Theorem 2.1.
Let $p\in[1,\infty)$. If $(v^{n})_{n\in\mathbb{N}}\subset L^{p}(\Omega)$ is such that
$\sup_{n\in\mathbb{N}}\|v^{n}\|_{p}<\infty\quad\text{and}\quad\lim_{h\to 0}\sup_{n\in\mathbb{N}}\|\tau_{h}\tilde{v}^{n}-\tilde{v}^{n}\|_{L^{p}(\mathbb{R}^{d})}=0,$
then, up to a subsequence, $v^{n}\to v$ in $L^{p}(\Omega)$ for some $v\in L^{p}(\Omega)$.
We also recall the following well-known characterization of Sobolev and $BV$ functions, see [Bre11book]*Prop. 9.3 and Rem. 6 for example.
Theorem 2.2.
Let $p\in[1,\infty)$ and $v\in L^{p}(\mathbb{R}^{d})$. The following are equivalent:
  1. (i)
    $v\in W^{1,p}(\mathbb{R}^{d})$ for $p>1$ or $v\in BV(\mathbb{R}^{d})$ for $p=1$;
  2. (ii)
    $\sup_{\|h\|\leq 1}\|\tau_{h}v-v\|_{p}<\infty$.
3. Proof of Theorem˜1.1
Throughout this section, we let $p\in[1,\infty)$, $(s_{n})_{n\in\mathbb{N}}\subset(0,1)$ be such that $s_{n}\to 1^{-}$, and $(f_{n})_{n\in\mathbb{N}}\subset L^{\infty}(\mathbb{R}^{d};[0,\infty))$ and $f\in\mathrm{Lip}_{b}(\mathbb{R}^{d};(0,\infty))$ be such that $f_{n}\to f$ in $L^{\infty}(\mathbb{R}^{d})$.
We preliminarily prove Theorem˜1.1 in the case $f_{n}=f$ for every $n\in\mathbb{N}$. We restate our result in this particular case for better clarity.
Theorem 3.1.
The following hold.
  1. (i)
    (Compactness) If $(u^{n})_{n\in\mathbb{N}}\subset L^{p}(\Omega)$ is such that
    $\sup_{n\in\mathbb{N}}\left((1-s_{n})[\tilde{u}^{n}]^{p}_{s_{n},p,f}+\|u^{n}\|_{L^{p}(\Omega)}^{p}\right)<\infty,$ (3.1)
    then, up to a subsequence, $u^{n}\to u$ in $L^{p}(\Omega)$ for some $u\in W^{1,p}_{0}(\Omega)$ if $p\in(1,\infty)$ or $u\in BV(\Omega)$ if $p=1$.
  2. (ii)
    ($\Gamma$-liminf inequality) If $(u_{n})_{n\in\mathbb{N}}\subset L^{p}(\Omega)$ is such that $u_{n}\to u$ in $L^{p}(\Omega)$ for some $u\in L^{p}(\Omega)$, then
    $\begin{split}&K_{d,p}\|\nabla u\|^{p}_{p,f^{2}}\leq\liminf_{n\to\infty}(1-s_{n})[\tilde{u}^{n}]^{p}_{s_{n},p,f}\quad\text{for $p\in(1,\infty)$},\\ &K_{d,1}\|Du\|_{1,f^{2}}\leq\liminf_{n\to\infty}(1-s_{n})[\tilde{u}^{n}]_{s_{n},1,f}\quad\text{for $p=1$}.\end{split}$ (3.2)
  3. (iii)
    ($\Gamma$-limsup inequality) If $u\in W^{1,p}_{0}(\Omega)$ if $p>1$, or $u\in BV(\Omega)$ if $p=1$, then there exists $(u^{n})_{n\in\mathbb{N}}\subset L^{p}(\Omega)$ such that $u_{n}\to u$ in $L^{p}(\Omega)$ and
    $\begin{split}&K_{d,p}\|\nabla u\|^{p}_{p,f^{2}}=\lim_{n\to\infty}(1-s_{n})[\tilde{u}^{n}]^{p}_{s_{n},p,f}\quad\text{for $p\in(1,\infty)$},\\ &K_{d,1}\|Du\|_{1,f^{2}}=\lim_{n\to\infty}(1-s_{n})[\tilde{u}^{n}]_{s_{n},1,f}\quad\text{for $p=1$.}\end{split}$ (3.3)
The proof of the three statements (i), (ii), and (iii) of Theorem˜3.1 is split across Sections˜3.1, 3.2 and 3.3. The proof of Theorem˜1.1 is given in Section˜3.4.
3.1. Proof of Theorem˜3.1(i)
We adapt the strategy of [AdPM01] to our setting. To this aim, we need two preliminary results. The first one is the following, which generalizes [AdPM01]*Prop. 5 to any $p\in[1,\infty)$ and weighted $L^{p}$ norms. We also refer to [CdLKNP23]*Prop. 2.4 for the case $p=2$ without weights.
Proposition 3.2.
Let $f\in\mathrm{Lip}_{b}(\mathbb{R}^{d};(0,\infty))$. There exists $C=C(d,p)>0$ such that
$\|\tau_{h}v-v\|_{L^{p}_{f}(E)}^{p}\leq C\frac{\|h\|^{p}}{\varrho^{d+p}}\int_{B_{\varrho}}\|\tau_{y}v-v\|_{L^{p}_{f}(E_{\|h\|})}^{p}\,\mathrm{d}y$ (3.4)
for every $v\in L^{p}(\mathbb{R}^{d})$, $h\in\mathbb{R}^{d}$, $\varrho\in(0,\|h\|]$, and every bounded open set $E\subset\mathbb{R}^{d}$, where $E_{\|h\|}$ is defined according to the notation in (2.1).
Proof.
The proof closely follows the one of [AdPM01]*Prop. 5. Let $\varphi\in C_{c}^{1}(B_{1})$ be such that
$\varphi\geq 0\quad\text{and}\quad\int_{B_{1}}\varphi(x)\,\mathrm{d}x=1.$ (3.5)
For every $\varrho>0$, we let $U_{\varrho}$ and $V_{\varrho}$ be defined as
$U_{\varrho}(x)$ $=\frac{1}{\varrho^{d}}\int_{B_{\varrho}}v(x+y)\varphi\left(\frac{y}{\varrho}\right)\,\mathrm{d}y,$
$V_{\varrho}(x)$ $=\frac{1}{\varrho^{d}}\int_{B_{\varrho}}(v(x)-v(x+y))\varphi\left(\frac{y}{\varrho}\right)\,\mathrm{d}y,$
for every $x\in\mathbb{R}^{d}$. Owing to (3.5), we have that $v(x)=U_{\varrho}(x)+V_{\varrho}(x)$ for every $\varrho>0$ and $x\in\mathbb{R}^{d}$, so that
$|\tau_{h}v(x)-v(x)|^{p}\leq 3^{p}\big(|U_{\varrho}(x+h)-U_{\varrho}(x)|^{p}+|V_{\varrho}(x)|^{p}+|V_{\varrho}(x+h)|^{p}\big).$ (3.6)
We now estimate each term in the right-hand side of (3.6) separately. Concerning the second and third term, by Jensen’s inequality, we can estimate
$|V_{\varrho}(\xi)|^{p}\leq\frac{\omega_{d}}{\varrho^{d}}\,\|\varphi\|^{p}_{\infty}\int_{B_{\varrho}}|v(\xi)-\tau_{y}v(\xi)|^{p}\,\mathrm{d}y,$ (3.7)
for every $\xi\in\mathbb{R}^{d}$. Instead, concerning the first term, by the change of variables $z=x+y$, we can rewrite
$U_{\varrho}(x)=\frac{1}{\varrho^{d}}\int_{B_{\varrho}(x)}v(z)\,\varphi\left(\frac{z-x}{\varrho}\right)\,\mathrm{d}z.$
Thus, owing to the fact that $\varphi((z-\cdot)\varrho^{-1})\in C^{1}_{c}(B_{\varrho}(x))$, we can integrate by parts and get
$\begin{split}\nabla U_{\varrho}(x)=&-\frac{1}{\varrho^{d+1}}\int_{B_{\varrho}(x)}v(z)\,\nabla\varphi\left(\frac{z-x}{\varrho}\right)\,\mathrm{d}z\\ =&-\frac{1}{\varrho^{d+1}}\int_{B_{\varrho}(x)}(v(z)-v(x))\,\nabla\varphi\left(\frac{z-x}{\varrho}\right)\,\mathrm{d}z\\ =&-\frac{1}{\varrho^{d+1}}\int_{B_{\varrho}}(v(x+y)-v(x))\,\nabla\varphi\left(\frac{y}{\varrho}\right)\,\mathrm{d}y.\end{split}$
Therefore, by the Fundamental Theorem of Calculus and by Jensen’s inequality, we obtain
$|U_{\varrho}(x$ $+h)-U_{\varrho}(x)|^{p}\leq\|h\|^{p}\int_{0}^{1}|\nabla U_{\varrho}(x+th)|^{p}\,\mathrm{d}t$
$\leq\omega_{d}^{p-1}\frac{\|h\|^{p}}{\varrho^{d+p}}\|\nabla\varphi\|_{\infty}^{p}\int_{0}^{1}\int_{B_{\varrho}}|\tau_{y}v(x+th)-v(x+th)|^{p}\,\mathrm{d}y\,\mathrm{d}t.$ (3.8)
Now, using that $\varrho<\|h\|$ and combining (3.6), (3.7), and (3.1), we get that
$|\tau_{h}v(x)-v(x)|^{p}$ $\leq C\frac{\|h\|^{p}}{\varrho^{d+p}}\int_{0}^{1}\int_{B_{\varrho}}|\tau_{y}v(x+th)-v(x+th)|^{p}\,\mathrm{d}y\,\mathrm{d}t$
$+C\frac{\|h\|^{p}}{\varrho^{d+p}}\int_{B_{\varrho}}|\tau_{y}v(x)-v(x)|^{p}\,\mathrm{d}y$
$+C\frac{\|h\|^{p}}{\varrho^{d+p}}\int_{B_{\varrho}}|\tau_{y}v(x+h)-v(x+h)|^{p}\,\mathrm{d}y,$ (3.9)
where we have set
$C=C(d,p)=(3\max\{\|\varphi\|_{\infty};\|\nabla\varphi\|_{\infty}\})^{p}\,\omega_{d}.$
Multiplying inequality (3.9) by $f(x)$ and integrating with respect to $x\in E$, the claim immediately follows by Fubini’s Theorem. We omit the simple details. ∎
We can now pass to the following result, which extends [AdPM01]*Prop. 4 to any $p\in[1,\infty)$ also in the case of weighted $L^{p}$ norms.
Proposition 3.3.
Let $f\in\mathrm{Lip}_{b}(\mathbb{R}^{d};(0,\infty))$. There exists $C=C(d,p)>0$ such that
$\|\tau_{h}v-v\|_{L^{p}_{f}(E)}^{p}\leq C(1-s)\|h\|^{sp}\int_{B_{\|h\|}}\frac{\|\tau_{y}v-v\|_{L^{p}_{f}(E_{\|h\|})}^{p}}{\|y\|^{d+sp}}\,\mathrm{d}y$
for every $v\in L^{p}(\mathbb{R}^{d})$, $h\in\mathbb{R}^{d}$, and every bounded open set $E\subset\mathbb{R}^{d}$, where $E_{\|h\|}$ is defined according to the notation in (2.1).
The proof of Proposition˜3.3 requires the following Hardy-type inequality, which is taken from [AdPM01]*Prop. 6.
Lemma 3.4.
If $\varphi\colon\mathbb{R}\to[0,\infty)$ is a Borel function, then
$\int_{0}^{r}\frac{1}{\varrho^{d+l+1}}\int_{0}^{\varrho}\varphi(t)\,\mathrm{d}t\,\mathrm{d}\varrho\leq\frac{1}{d+l}\int_{0}^{r}\frac{\varphi(t)}{t^{d+l}}\,\mathrm{d}t,$ (3.10)
for every $l,r\geq 0$.
Actually, in the proof of Proposition˜3.3 we use the weaker estimate
$\int_{0}^{r}\frac{1}{\varrho^{d+l+1}}\int_{0}^{\varrho}\varphi(t)\,\mathrm{d}t\,\mathrm{d}\varrho\leq\frac{1}{d}\int_{0}^{r}\frac{\varphi(t)}{t^{d+l}}\,\mathrm{d}t,$ (3.11)
that is, we can ignore the dependence on $l$ in the prefactor in the right-hand side of (3.10).
Proof of Proposition˜3.3.
We let $\varphi_{v}\colon[0,\|h\|]\to\mathbb{R}$ be defined as
$\varphi_{v}(t)=\int_{\partial B_{t}}\|\tau_{y}v-v\|_{L^{p}_{f}(E_{\|h\|})}^{p}\,\mathrm{d}\mathcal{H}^{d-1}(y),$ (3.12)
for all $t>0$. Owing to (3.4) and to the definition in (3.12), we can estimate
$\|\tau_{h}v-v\|_{L_{f}^{p}(E)}^{p}\leq C\frac{\|h\|^{p}}{\varrho^{d+p}}\int_{0}^{\varrho}\varphi_{v}(t)\,\mathrm{d}t,$ (3.13)
for some $C=C(d,p)>0$. We now multiply both sides of (3.13) by $\varrho^{-1+p-sp}$ and integrate in the interval $0,\|h\|]$ with respect to $\varrho$, getting
$\|\tau_{h}v-v\|_{L_{f}^{p}(E)}^{p}\leq Cp(1-s)\frac{\|h\|^{p}}{\|h\|^{p-sp}}\int_{0}^{\|h\|}\frac{1}{\varrho^{d+sp+1}}\int_{0}^{\varrho}\varphi_{v}(t)\,\mathrm{d}t\,\mathrm{d}\varrho.$
By exploiting (3.11) with $l=sp$ and $\varphi=\varphi_{v}$, we thus obtain that
$\|\tau_{h}v-v\|_{L^{p}_{f}(E)}^{p}\leq C(1-s)\|h\|^{sp}\int_{0}^{\|h\|}\frac{\varphi_{v}(t)}{t^{d+sp}}\,\mathrm{d}t,$
and the conclusion follows from the very definition of $\varphi_{v}$. ∎
We are now ready to detail the proof of the compactness statement (i) in Theorem˜3.1.
Proof of Theorem˜3.1(i).
Given $h\in\mathbb{R}^{d}$ such that $\|h\|<1$, we have $\Omega_{\|h\|}\Subset\Omega_{1}\Subset(\Omega_{1})_{\|h\|}$ (recall the notation in (2.1)). We can hence set
$c=c(\Omega,f)=\inf_{(\Omega_{1})_{\|h\|}}f>0,$
and observe that
$c^{2}\|\tau_{h}\tilde{u}^{n}-\tilde{u}^{n}\|^{p}_{L^{p}(\mathbb{R}^{d})}=c^{2}\|\tau_{h}\tilde{u}^{n}-\tilde{u}^{n}\|^{p}_{L^{p}(\Omega_{\|h\|})}\leq c\|\tau_{h}\tilde{u}^{n}-\tilde{u}^{n}\|^{p}_{L^{p}_{f}(\Omega_{1})}.$
By Proposition˜3.3 applied on $\Omega_{1}$ and by the previous inequality, we have
$c^{2}\|\tau_{h}\tilde{u}^{n}-\tilde{u}^{n}\|^{p}_{L^{p}(\mathbb{R}^{d})}\leq C(1-s_{n})\|h\|^{s_{n}p}\int_{B_{\|h\|}}\frac{c\|\tau_{y}\tilde{u}^{n}-\tilde{u}^{n}\|_{L^{p}_{f}((\Omega_{1})_{\|h\|})}^{p}}{\|y\|^{d+s_{n}p}}\,\mathrm{d}y,$
where $C=C(d,p)>0$. Now, explicitly writing down the $L^{p}_{f}$ norm on the right-hand side, swapping order of integration, performing the change of variables $y=\xi-x$, and bounding $c$ with $f(\xi)$, we obtain that
$c^{2}\|\tau_{h}\tilde{u}^{n}$ $-\tilde{u}^{n}\|^{p}_{L^{p}(\mathbb{R}^{d})}\leq C(1-s_{n})\|h\|^{s_{n}p}\int_{B_{\|h\|}}\int_{(\Omega_{1})_{\|h\|}}\frac{|\tilde{u}^{n}(x+y)-\tilde{u}^{n}(x)|^{p}}{\|y\|^{d+s_{n}p}}cf(x)\,\mathrm{d}x\,\mathrm{d}y$
$\leq C(1-s_{n})\|h\|^{s_{n}p}\int_{(\Omega_{1})_{\|h\|}}\int_{B_{\|h\|}(x)}\frac{|\tilde{u}^{n}(\xi)-\tilde{u}^{n}(x)|^{p}}{\|\xi-x\|^{d+s_{n}p}}f(\xi)f(x)\,\mathrm{d}\xi\,\mathrm{d}x$
$\leq C(1-s_{n})\|h\|^{s_{n}p}[\tilde{u}^{n}]^{p}_{s_{n},p,f}.$
Dividing by $c^{2}$ and owing to our equiboundedness assumption (3.1), we get that
$\|\tau_{h}\tilde{u}^{n}-\tilde{u}^{n}\|_{L^{p}(\mathbb{R}^{d})}\leq C(d,p,M,\Omega,f)\|h\|^{s_{n}},$ (3.14)
for all $h\in\mathbb{R}^{d}$ such that $\|h\|\leq 1$. Thus, owing to (3.1) and to (3.14), we can apply Theorem˜2.1 and find $u\in L^{p}(\mathbb{R}^{d})$ such that, up to a subsequence, $\tilde{u}^{n}\to u$ in $L^{p}(\mathbb{R}^{d})$. Furthermore, since $\tilde{u}^{n}=0$ for all $n\in\mathbb{N}$ on $\mathbb{R}^{d}\setminus\Omega$, we also have that $u=0$ on $\mathbb{R}^{d}\setminus\Omega$. Finally, letting $n\to\infty$ in (3.14), we have
$\|\tau_{h}u-u\|_{L^{p}(\mathbb{R}^{d})}\leq C(d,p,M,\Omega,f)\|h\|,$
for all $h$ with $\|h\|\leq 1$, so that $u\in W^{1,p}(\mathbb{R}^{d})$ for $p>1$ or $u\in BV(\mathbb{R}^{d})$ for $p=1$ by Theorem˜2.2. Since $u=0$ on $\mathbb{R}^{d}\setminus\Omega$ and $\Omega$ has Lipschitz boundary, we get that $u|_{\Omega}\in W^{1,p}_{0}(\Omega)$ for $p>1$, or $u|_{\Omega}\in BV(\Omega)$ for $p=1$, concluding the proof. ∎
3.2. Proof of Theorem˜3.1(ii)
We adapt the strategy of the proof of [CdLKNP23]*Thm. 2.1 to our setting. To this aim, we need some preliminaries.
Let us begin with some notation. We let $Q=(-1,1)^{d}$. Consequently, given $\gamma>0$, for every $i\in\gamma\mathbb{Z}^{d}$ and $a\geq 0$, we let $Q_{i}^{a}=i+aQ$. Note that, if $a=\gamma$, then the family of cubes $(Q^{\gamma}_{i})_{i}$ is a tiling of $\mathbb{R}^{d}$. Moreover, since $\Omega$ is bounded, the set
[…]
is finite. In addition, given $f\colon\mathbb{R}^{d}\to\mathbb{R}$, we let $f_{i}^{a}=\inf_{Q_{i}^{a}}f$. Notice that, whenever $f\in\mathrm{Lip}(\mathbb{R}^{d};(0,\infty))$, then $f_{i}^{a}>0$. Finally, we let $\eta\in C_{c}^{\infty}(B_{1})$ be such that $\eta\geq 0$ and $\int_{B_{1}}\eta\,\mathrm{d}x=1$ and, for every $\varepsilon>0$, we set $\eta_{\varepsilon}(\,\cdot\,)=\varepsilon^{-d}\eta(\,\cdot\,/\varepsilon)$. Accordingly, we let $u_{\varepsilon}=u\ast\eta_{\varepsilon}$ for every $\varepsilon>0$ and $u\in L^{1}_{\rm loc}(\mathbb{R}^{d})$.
We can now prove the following preliminary estimate.
Lemma 3.5.
Let $p\in[1,\infty)$, $f\in\mathrm{Lip}_{b}(\mathbb{R}^{d};(0,\infty))$. There exist $\varepsilon,\beta,\gamma>0$ with $\varepsilon\ll\beta\ll\gamma$ such that
$\begin{split}\int_{Q_{i}^{(1-\beta)\gamma}}\int_{Q_{i}^{(1-\beta)\gamma}}&\frac{|\tilde{u}_{\varepsilon}(x)-\tilde{u}_{\varepsilon}(y)|^{p}}{\|x-y\|^{d+sp}}(f_{i}^{\gamma})^{2}\,\mathrm{d}x\,\mathrm{d}y\\ &\quad\qquad\leq\int_{Q_{i}^{\gamma}}\int_{Q_{i}^{\gamma}}\frac{|\tilde{u}(x)-\tilde{u}(y)|^{p}}{\|x-y\|^{d+sp}}f(x)f(y)\,\mathrm{d}x\,\mathrm{d}y\end{split}$ (3.15)
holds for all $i\in\gamma\mathbb{Z}^{d}$ and $u\in L^{p}(\Omega)$.
Proof.
Clearly, we can choose $\varepsilon$, $\beta$, and $\gamma$ such that
$Q_{i}^{(1-\beta)\gamma}+z\subset Q_{i}^{\gamma}$ (3.16)
for all $z\in B_{\varepsilon}$ and all $i\in\gamma\mathbb{Z}^{d}$. Indeed,
[…]
and therefore, to get (3.16), it is enough to choose $\varepsilon<\beta\gamma$ to ensure $(1-\beta)\gamma+\varepsilon<\gamma$. Using the definition of convolution, Jensen’s inequality, changing order of integration, performing the change of variables $x-z=\xi$ and $y-z=\zeta$, owing to (3.16), changing once again order of integration, using the fact that $\eta_{\varepsilon}$ has unit $L^{1}$ norm, and finally owing to the definition of $f_{i}^{\gamma}=\inf_{Q_{i}^{\gamma}}f$, we obtain
$\int_{Q_{i}^{(1-\beta)\gamma}}$ $\int_{Q_{i}^{(1-\beta)\gamma}}\frac{|\tilde{u}_{\varepsilon}(x)-\tilde{u}_{\varepsilon}(y)|^{p}}{\|x-y\|^{d+sp}}(f_{i}^{\gamma})^{2}\,\mathrm{d}x\,\mathrm{d}y$
$\leq\int_{Q_{i}^{(1-\beta)\gamma}}\int_{Q_{i}^{(1-\beta)\gamma}}\int_{B_{\varepsilon}}\frac{|\tilde{u}(x-z)-\tilde{u}(y-z)|^{p}}{\|x-z-(y-z)\|^{d+sp}}\eta_{\varepsilon}(z)(f_{i}^{\gamma})^{2}\,\mathrm{d}z\,\mathrm{d}x\,\mathrm{d}y$
$\leq\int_{B_{\varepsilon}}\int_{Q_{i}^{(1-\beta)\gamma}+z}\int_{Q_{i}^{(1-\beta)\gamma}+z}\frac{|\tilde{u}(\xi)-\tilde{u}(\zeta)|^{p}}{\|\xi-\zeta\|^{d+sp}}\eta_{\varepsilon}(z)(f_{i}^{\gamma})^{2}\,\mathrm{d}\xi\,\mathrm{d}\zeta\,\mathrm{d}z$
$\leq\int_{Q_{i}^{\gamma}}\int_{Q_{i}^{\gamma}}\frac{|\tilde{u}(\xi)-\tilde{u}(\zeta)|^{p}}{\|\xi-\zeta\|^{d+sp}}\,f(\xi)\,f(\zeta)\,\mathrm{d}\xi\,\mathrm{d}\zeta,$
concluding the proof. ∎
We are now ready to prove the $\liminf$ statement (ii) in Theorem˜3.1.
Proof of Theorem˜3.1(ii).
Let $(u^{n})_{n\in\mathbb{N}}\subset L^{p}(\Omega)$ be such that $u^{n}\to u$ in $L^{p}(\Omega)$ for some $u\in L^{p}(\Omega)$. As a consequence, $\sup_{n\in\mathbb{N}}\|u^{n}\|_{p}<\infty$. Furthermore, we can also assume that
$\liminf_{n\to\infty}(1-s_{n})[\tilde{u}^{n}]^{p}_{s_{n},p,f}<\infty,$
otherwise inequality (3.2) is trivially true. Therefore, we can assume the validity of (3.1), which, in turn, implies that $u\in W^{1,p}_{0}(\Omega)$ for $p>1$, or $u\in BV(\Omega)$ for $p=1$. By Lemma˜3.5, there exist $\varepsilon,\beta,\gamma>0$ with $\varepsilon\ll\beta\ll\gamma$ such that
$\begin{split}\int_{Q_{i}^{\gamma}}\int_{Q_{i}^{\gamma}}&\frac{|\tilde{u}^{n}(x)-\tilde{u}^{n}(y)|^{p}}{\|x-y\|^{d+s_{n}p}}f(x)f(y)\,\mathrm{d}y\,\mathrm{d}x\\ &\quad\geq\int_{Q_{i}^{(1-\beta)\gamma}}\int_{Q_{i}^{(1-\beta)\gamma}}\frac{|(\tilde{u}^{n})_{\varepsilon}(x)-(\tilde{u}^{n})_{\varepsilon}(y)|^{p}}{\|x-y\|^{d+s_{n}p}}(f_{i}^{\gamma})^{2}\,\mathrm{d}y\,\mathrm{d}x,\end{split}$ (3.17)
for every $i\in\gamma\mathbb{Z}^{d}$ and $n\in\mathbb{N}$. We now perform a first-order Taylor expansion of $(\tilde{u}^{n})_{\varepsilon}$. Precisely, owing to the uniform bound on the $p$-norms granted by (3.1) and the boundedness of $\Omega$, we can estimate
$|\nabla(\tilde{u}^{n})_{\varepsilon}(x)\cdot(x-y)|$ $\leq|(\tilde{u}^{n})_{\varepsilon}(x)-(\tilde{u}^{n})_{\varepsilon}(y)|+\frac{1}{2}|\langle D^{2}(\tilde{u}^{n})_{\varepsilon}(\xi)(x-y),(x-y)\rangle|$
$\leq|(\tilde{u}^{n})_{\varepsilon}(x)-(\tilde{u}^{n})_{\varepsilon}(y)|+C(\Omega)\|u^{n}\|_{1}\,\|\eta_{\varepsilon}\|_{C^{2}(\mathbb{R}^{d})}\,\|x-y\|^{2}$
$\leq|(\tilde{u}^{n})_{\varepsilon}(x)-(\tilde{u}^{n})_{\varepsilon}(y)|+C(\varepsilon,M,\Omega)\|x-y\|^{2},$ (3.18)
where $\xi$ belongs to the segment from $x$ to $y$. Now, assuming $\|x-y\|$ small enough (which is always possible by taking $\gamma$ small enough), since $\tilde{u}^{n}_{\varepsilon}$ is locally Lipschitz, we have
$|(\tilde{u}^{n})_{\varepsilon}(x)-(\tilde{u}^{n})_{\varepsilon}(y)|$ $+C(\varepsilon,M,\Omega)\|x-y\|^{2}$
$\leq\|\nabla(\tilde{u}^{n})_{\varepsilon}\|_{\infty}\|x-y\|+C(\varepsilon,M,\Omega)\|x-y\|^{2}$
$=C(\Omega)\|u^{n}\|_{1}\|\eta_{\varepsilon}\|_{C^{1}(\mathbb{R}^{d})}\|x-y\|+C(\varepsilon,M,\Omega)\|x-y\|^{2}$
$=C(\varepsilon,M,\Omega)\|x-y\|.$ (3.19)
Taking $p$-th powers in (3.18), using that $(t_{0}+t)^{p}=t_{0}^{p}+p(t_{0}+\tau)^{p-1}t$ with $\tau\in(0,t)$, and finally owing to (3.2), we get
$\left|\nabla(\tilde{u}^{n})_{\varepsilon}(x)\cdot(x-y)\right|^{p}\leq|(\tilde{u}^{n})_{\varepsilon}(x)-(\tilde{u}^{n})_{\varepsilon}(y)|^{p}+C(p,\varepsilon,M,\Omega)\|x-y\|^{p+1}.$
Therefore, plugging the above inequality in the inner integral in the right-hand side of (3.17), we get that
$\int_{Q_{i}^{(1-\beta)\gamma}}\frac{|(\tilde{u}^{n})_{\varepsilon}(x)-(\tilde{u}^{n})_{\varepsilon}(y)|^{p}}{\|x-y\|^{d+s_{n}p}}(f_{i}^{\gamma})^{2}\,\mathrm{d}y\geq I^{\prime}+I^{\prime\prime},$ (3.20)
having set
$\begin{split}I^{\prime}&=\int_{Q_{i}^{(1-\beta)\gamma}}\frac{|\nabla(\tilde{u}^{n})_{\varepsilon}(x)\cdot(x-y)|^{p}}{\|x-y\|^{d+s_{n}p}}(f_{i}^{\gamma})^{2}\,\mathrm{d}y,\\ I^{\prime\prime}&=-C\int_{Q_{i}^{(1-\beta)\gamma}}\frac{\|x-y\|^{p+1}}{\|x-y\|^{d+s_{n}p}}(f_{i}^{\gamma})^{2}\,\mathrm{d}y.\end{split}$ (3.21)
We now estimate the two terms in (3.21) separately. On the one hand, we have
$I^{\prime\prime}$ $\geq-C(p,\varepsilon,M,\Omega)\|f\|^{2}_{\infty}\int_{B_{\delta_{1}}}\frac{\mathrm{d}\xi}{\|\xi\|^{d+s_{n}p-p-1}}$
$=-C(p,\varepsilon,M,\Omega,f)d\omega_{d}\int_{0}^{\delta_{1}}\frac{\mathrm{d}\varrho}{\varrho^{s_{n}p-p}}=-\frac{C(d,p,\varepsilon,M,\Omega,f)}{1+p-s_{n}p}\delta_{1}^{1+p-s_{n}p}$
for $x\in Q_{i}^{(1-\beta)\gamma}$, where we have set $\delta_{1}=2\sqrt{d}(1-\beta)\gamma$. Integrating the above inequality over the cube $Q_{i}^{(1-\beta)\gamma}$ with respect to $x$, we obtain
$\int_{Q_{i}^{(1-\beta)\gamma}}I^{\prime\prime}\,\mathrm{d}x\geq-\frac{C(d,p,\varepsilon,M,\Omega,f)}{1+p-s_{n}p}\big|Q_{i}^{(1-\beta)\gamma}\big|\delta_{1}^{1+p-s_{n}p},$
so that
$\liminf_{n\to\infty}(1-s_{n})\int_{Q_{i}^{(1-\beta)\gamma}}I^{\prime\prime}\,\mathrm{d}x\geq 0.$ (3.22)
On the other hand, calling $\delta_{2}=\delta_{2}(x)=\operatorname{dist}\big(x;\partial Q_{i}^{(1-\beta)\gamma}\big)$, taking the normalization $\nu(x)=\nabla(\tilde{u}^{n})_{\varepsilon}(x)/\|\nabla(\tilde{u}^{n})_{\varepsilon}(x)\|$, applying the change of variables z=xyz=x-y, exploiting the invariance of the integral on $B_{\delta_{2}}$ with respect to rotations, applying the change of variable $z\mapsto z\delta_{2}^{-1}$, and observing that, for every $\nu\in\mathbb{S}^{d-1}$
$(1-s_{n})\int_{B_{1}}\frac{|z\cdot\nu|^{p}}{\|z\|^{d+s_{n}p}}\,\mathrm{d}z=(1-s_{n})\int_{0}^{1}\varrho^{p-s_{n}p-1}\,\int_{\partial B_{1}}|\eta\cdot\nu|^{p}\,\mathrm{d}\eta\,\mathrm{d}\varrho=K_{d,p},$ (3.23)
we get that
$I^{\prime}$ $=|\nabla(\tilde{u}^{n})_{\varepsilon}(x)|^{p}(f_{i}^{\gamma})^{2}\int_{Q_{i}^{(1-\beta)\gamma}}\frac{|\nu(x)\cdot(x-y)|^{p}}{\|x-y\|^{d+s_{n}p}}\,\mathrm{d}y$
$\geq|\nabla(\tilde{u}^{n})_{\varepsilon}(x)|^{p}(f_{i}^{\gamma})^{2}\int_{B_{\delta_{2}}}\frac{|\nu(x)\cdot z|^{p}}{\|z\|^{d+s_{n}p}}\,\mathrm{d}z=\frac{K_{d,p}|\nabla(\tilde{u}^{n})_{\varepsilon}(x)|^{p}\delta_{2}^{p(1-s_{n})}(f_{i}^{\gamma})^{2}}{1-s_{n}}.$
Multiplying the above inequality by $(1-s_{n})$, integrating over the cube $Q_{i}^{(1-\beta)\gamma}$ with respect to $x$, taking the $\liminf$ as $n\to\infty$, owing to Fatou’s lemma, we get that
$\liminf_{n\to\infty}(1-s_{n})\int_{Q_{i}^{(1-\beta)\gamma}}I^{\prime}\,\mathrm{d}x\geq K_{d,p}\int_{Q_{i}^{(1-\beta)\gamma}}\liminf_{n\to\infty}\big(|\nabla(\tilde{u}^{n})_{\varepsilon}(x)|^{p}\delta_{2}^{p(1-s_{n})}\big)(f_{i}^{\gamma})^{2}\,\mathrm{d}x.$
Since $(u^{n})_{\varepsilon}\to u_{\varepsilon}$ in any Sobolev norm as $n\to\infty$, up to passing to a suitable subsequence, $\nabla(\tilde{u}^{n})_{\varepsilon}(x)\to\nabla\tilde{u}_{\varepsilon}(x)$ for a.e. $x\in\mathbb{R}^{d}$. Therefore, we get that
$\liminf_{n\to\infty}(1-s_{n})\int_{Q_{i}^{(1-\beta)\gamma}}I^{\prime}\,\mathrm{d}x\geq K_{d,p}\int_{Q_{i}^{(1-\beta)\gamma}}|\nabla\tilde{u}_{\varepsilon}(x)|^{p}(f_{i}^{\gamma})^{2}\,\mathrm{d}x.$ (3.24)
Thence, by taking the $\liminf$ as $n\to\infty$ of $(1-s_{n})$ times (3.17) and owing to (3.20)–(3.24), we have
$\begin{split}\liminf_{n\to\infty}(1-s_{n})\int_{Q_{i}^{\gamma}}\int_{Q_{i}^{\gamma}}\frac{|\tilde{u}^{n}(x)-\tilde{u}^{n}(y)|^{p}}{\|x-y\|^{d+sp}}&f(x)f(y)\,\mathrm{d}y\,\mathrm{d}x\\ &\geq K_{d,p}\int_{Q_{i}^{(1-\beta)\gamma}}|\nabla\tilde{u}_{\varepsilon}(x)|^{p}(f_{i}^{\gamma})^{2}\,\mathrm{d}x\end{split}$
whenever $\varepsilon\ll\beta\ll\gamma\ll 1$. Now first letting $\varepsilon\to 0^{+}$, and then $\beta\to 0^{+}$, we get
$\begin{split}\liminf_{n\to\infty}(1-s_{n})\int_{Q_{i}^{\gamma}}\int_{Q_{i}^{\gamma}}&\frac{|\tilde{u}^{n}(x)-\tilde{u}^{n}(y)|^{p}}{\|x-y\|^{d+sp}}f(x)f(y)\,\mathrm{d}y\,\mathrm{d}x\\ &\geq K_{d,p}\int_{Q_{i}^{\gamma}}|\nabla\tilde{u}(x)|^{p}(f_{i}^{\gamma})^{2}\,\mathrm{d}x,\quad\text{ if $p\in(1,\infty)$},\\ \liminf_{n\to\infty}(1-s_{n})\int_{Q_{i}^{\gamma}}\int_{Q_{i}^{\gamma}}&\frac{|\tilde{u}^{n}(x)-\tilde{u}^{n}(y)|}{\|x-y\|^{d+s}}f(x)f(y)\,\mathrm{d}y\,\mathrm{d}x\\ &\geq K_{d,1}\int_{Q_{i}^{\gamma}}(f_{i}^{\gamma})^{2}\,\mathrm{d}|Du|(x),\quad\text{ if $p=1$}.\end{split}$ (3.25)
We notice now that, by the Lebesgue’s Dominated Convergence Theorem,
$\begin{split}\lim_{\gamma\to 0^{+}}\sum_{i\in I_{\gamma}}K_{d,p}\int_{Q_{i}^{\gamma}}|\nabla\tilde{u}(x)|^{p}(f_{i}^{\gamma})^{2}\,\mathrm{d}x=K_{d,p}\|\nabla u\|^{p}_{p,f^{2}}\\ \text{and}\quad\lim_{\gamma\to 0^{+}}\sum_{i\in I_{\gamma}}K_{d,1}\int_{Q_{i}^{\gamma}}(f_{i}^{\gamma})^{2}\,\mathrm{d}|Du|=K_{d,1}\|Du\|_{1,f^{2}}.\end{split}$
Pairing this with (3.25) allows to conclude. Indeed, if $p\in(1,\infty)$ (the case $p=1$ being analogous and thus omitted), then we have
$K_{d,p}$ $\int_{\Omega}|\nabla u(x)|^{p}f^{2}(x)\,\mathrm{d}x$
$\leq\lim_{\gamma\to 0^{+}}\sum_{i\in I_{\gamma}}\liminf_{n\to\infty}(1-s_{n})\int_{Q_{i}^{\gamma}}\int_{Q_{i}^{\gamma}}\frac{|\tilde{u}^{n}(x)-\tilde{u}^{n}(y)|^{p}}{\|x-y\|^{d+s_{n}p}}f(x)f(y)\,\mathrm{d}y\,\mathrm{d}x$
$=\lim_{\gamma\to 0^{+}}\liminf_{n\to\infty}\sum_{i\in I_{\gamma}}(1-s_{n})\int_{Q_{i}^{\gamma}}\int_{Q_{i}^{\gamma}}\frac{|\tilde{u}^{n}(x)-\tilde{u}^{n}(y)|^{p}}{\|x-y\|^{d+s_{n}p}}f(x)f(y)\,\mathrm{d}y\,\mathrm{d}x$
$\leq\lim_{\gamma\to 0^{+}}\liminf_{n\to\infty}(1-s_{n})\int_{\Omega}\int_{\Omega}\frac{|\tilde{u}^{n}(x)-\tilde{u}^{n}(y)|^{p}}{\|x-y\|^{d+s_{n}p}}f(x)f(y)\,\mathrm{d}y\,\mathrm{d}x$
$=\liminf_{n\to\infty}(1-s_{n})\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|\tilde{u}^{n}(x)-\tilde{u}^{n}(y)|^{p}}{\|x-y\|^{d+s_{n}p}}f(x)f(y)\,\mathrm{d}y\,\mathrm{d}x,$
which is exactly the claimed inequality. ∎
3.3. Proof of Theorem˜3.1(iii)
We begin with the following preliminary result, establishing the $\limsup$ inequality (iii) in Theorem˜3.1 for smooth functions supported in $\Omega$.
Theorem 3.6.
If $v\in C_{c}^{\infty}(\mathbb{R}^{d})$, then
$\begin{split}\lim_{s\to 1^{-}}(1-s)\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|v(x)-v(y)|^{p}}{\|x-y\|^{d+sp}}&f(x)f(y)\,\mathrm{d}x\,\mathrm{d}y\\ &=K_{d,p}\int_{\mathbb{R}^{d}}\|\nabla v(z)\|^{p}f^{2}(z)\,\mathrm{d}z.\end{split}$ (3.26)
Proof.
Let us write
$\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|v(x)-v(y)|^{p}}{\|x-y\|^{d+sp}}f(x)f(y)\,\mathrm{d}x\,\mathrm{d}y=I_{1}+I_{2}+I_{3},$ (3.27)
where
$\begin{split}I_{1}&=\int_{\mathbb{R}^{d}}\int_{\{\|x-y\|<1\}}\frac{|v(x)-v(y)|^{p}}{\|x-y\|^{d+sp}}f^{2}(y)\,\mathrm{d}x\,\mathrm{d}y,\\ I_{2}&=\int_{\mathbb{R}^{d}}\int_{\{\|x-y\|<1\}}\frac{|v(x)-v(y)|^{p}}{\|x-y\|^{d+sp}}f(y)(f(x)-f(y))\,\mathrm{d}x\,\mathrm{d}y,\\ I_{3}&=\int_{\mathbb{R}^{d}}\int_{\{\|x-y\|\geq 1\}}\frac{|v(x)-v(y)|^{p}}{\|x-y\|^{d+sp}}f(x)f(y)\,\mathrm{d}x\,\mathrm{d}y.\end{split}$ (3.28)
We estimate the three terms separately. Concerning $I_{3}$, we have that
$I_{3}$ $\leq 2^{p}\|f\|^{2}_{\infty}\int_{\mathbb{R}^{d}}\int_{\{\|x-y\|\geq 1\}}\frac{|v(x)|^{p}+|v(y)|^{p}}{\|x-y\|^{d+sp}}\,\mathrm{d}x\,\mathrm{d}y$
$\leq 2^{p+1}\|f\|^{2}_{\infty}\int_{\mathbb{R}^{d}}\int_{B_{1}(y)^{c}}\frac{|v(y)|^{p}}{\|x-y\|^{d+sp}}\,\mathrm{d}x\,\mathrm{d}y$
$=2^{p+1}\|f\|^{2}_{\infty}\|v\|^{p}_{p}\int_{1}^{\infty}\frac{d\omega_{d}}{\varrho^{1+sp}}\,\mathrm{d}\varrho=\frac{2^{p+1}}{sp}\|f\|^{2}_{\infty}\|v\|^{p}_{p}.$ (3.29)
For $I_{2}$ instead, we can write, recalling the notation in (2.1),
$I_{2}$ $\leq\|f\|_{\infty}\|\nabla f\|_{\infty}\int_{\mathbb{R}^{d}}\int_{B_{1}(y)}\frac{|v(x)-v(y)|^{p}}{\|x-y\|^{d+sp-1}}\,\mathrm{d}x\,\mathrm{d}y$
$\leq C(f)\|\nabla v\|_{\infty}^{p}\int_{(\operatorname{spt}{v})_{1}}\int_{B_{1}(y)}\frac{1}{\|x-y\|^{d-1-p+sp}}\,\mathrm{d}x\,\mathrm{d}y$
$\leq C(f,v)d\omega_{d}|(\operatorname{spt}{v})_{1}|\int_{0}^{1}\frac{1}{\varrho^{-p+sp}}\,\mathrm{d}\varrho$
$\leq\frac{C(d,f,v)}{p-sp+1}\leq C(d,p,f,v).$ (3.30)
We are thus left with estimating $I_{1}$. To this aim, let us observe that
$|v(x)-v(y)|^{p}\leq|\nabla v(y)\cdot(x-y)|^{p}+\|D^{2}v\|_{\infty}^{2p}\,\|x-y\|^{2p}.$
Thus $I_{1}\leq I^{\prime}_{1}+I^{\prime\prime}_{1}$, where, owing to the fact that $v$ has compact support and recalling the notation in (2.1),
$\begin{split}I^{\prime}_{1}&=\int_{\mathbb{R}^{d}}\int_{B_{1}(y)}\frac{|\nabla v(y)\cdot(x-y)|^{p}}{\|x-y\|^{d+sp}}f^{2}(y)\,\mathrm{d}x\,\mathrm{d}y,\\ I^{\prime\prime}_{1}&=\int_{(\operatorname{spt}{v})_{1}}\int_{B_{1}(y)}\frac{\|D^{2}v\|_{\infty}^{2p}}{\|x-y\|^{d+sp-2p}}f^{2}(y)\,\mathrm{d}x\,\mathrm{d}y.\end{split}$
Now, on the one hand, we have that
$I^{\prime\prime}_{1}$ $\leq\|f\|_{\infty}^{2}\|D^{2}v\|_{\infty}^{2p}\int_{(\operatorname{spt}{v})_{1}}\int_{B_{1}(y)}\frac{1}{\|x-y\|^{d+sp-2p}}\,\mathrm{d}x\,\mathrm{d}y$
$\leq C(f,v)d\omega_{d}|(\operatorname{spt}{v})_{1}|\int_{0}^{1}\varrho^{p-1+p(1-s)}\,\mathrm{d}\varrho$
$=\frac{C(d,f,v)}{2p-sp}\leq C(d,p,f,v).$ (3.31)
By the non-negativity of $I^{\prime\prime}_{1},I_{2},I_{3}$, and by (3.29)–(3.31), we thus get that
$\lim_{s\to 1^{-}}(1-s)(I^{\prime\prime}_{1}+I_{2}+I_{3})=0.$ (3.32)
Therefore, in order to conclude, we are left with showing that $\lim_{s\to 1^{-}}(1-s)I^{\prime}_{1}$ equals the right hand side of (3.26). Indeed, by the change of variables z=xyz=x-y, by setting $\nu(y)=\nabla v(y)/\|\nabla v(y)\|$, and using (3.23), we have
${I^{\prime}}_{1}$ $=\int_{\mathbb{R}^{d}}\int_{B_{1}}\frac{|\nabla v(y)\cdot z|^{p}}{\|z\|^{d+sp}}f^{2}(y)\,\mathrm{d}z\,\mathrm{d}y$
$=\int_{\mathbb{R}^{d}}\|\nabla v(y)\|^{p}f^{2}(y)\int_{B_{1}}\frac{|\nu(y)\cdot z|^{p}}{\|z\|^{d+sp}}\,\mathrm{d}z\,\mathrm{d}y$
$=\frac{K_{d,p}}{(1-s)}\int_{\mathbb{R}^{d}}\|\nabla v(x)\|^{p}f^{2}(y)\,\mathrm{d}y.$ (3.33)
The conclusion hence follows by combining (3.32) and (3.33). ∎
Remark 3.7.
We underline that, in the chain of inequalities leading to (3.30), the Lipschitz regularity of the weight $f$ is crucial in order to ensure the finiteness of the integral in $\varrho$. If $f$ were only $\alpha$-Hölder, for some $\alpha\in(0,1)$, one would end up with $\varrho$ to the power $(\alpha-1-p+sp)$, which is not integrable in a neighborhood of the origin.
We are now ready to prove the $\limsup$ statement (iii) in Theorem˜3.1.
Proof of Theorem˜3.1(iii).
Let $u\in W^{1,p}_{0}(\Omega)$ for $p>1$, or $u\in BV(\Omega)$ for $p=1$. By the density of $C_{c}^{\infty}(\Omega)$ in $W^{1,p}_{0}(\Omega)$ for $p>1$, or in $BV(\Omega)$ for $p=1$, we can find $(v^{k})_{k\in\mathbb{N}}\subset C_{c}^{\infty}(\Omega)$ such that $v^{k}\to u$ in $W^{1,p}_{0}(\Omega)$ for $p>1$, or in energy in $BV(\Omega)$ for $p=1$. In view of Theorem˜3.6, we thus get that
$\lim_{n\to\infty}(1-s_{n})[\tilde{v}^{k}]^{p}_{s_{n},p,f}$ $=\lim_{n\to\infty}(1-s_{n})\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|\tilde{v}^{k}(x)-\tilde{v}^{k}(y)|^{p}}{\|x-y\|^{d+s_{n}p}}f(x)f(y)\,\mathrm{d}x\,\mathrm{d}y$
$=K_{d,p}\int_{\mathbb{R}^{d}}\|\nabla\tilde{v}^{k}(x)\|^{2}f^{2}(x)\,\mathrm{d}x$
$=K_{d,p}\int_{\Omega}\|\nabla v^{k}(x)\|^{p}f^{2}(x)\,\mathrm{d}x$
for every $k\in\mathbb{N}$. The conclusion hence follows by a standard diagonal argument. ∎
3.4. Proof of Theorem˜1.1
We can now prove Theorem˜1.1 in full generality. Indeed, the result easily follows by combining Theorem˜3.1 with Lemmas˜3.8 and 3.9 below (under the same standing assumptions as stated at the very beginning of Section˜3).
Lemma 3.8.
If $(u^{n})_{n\in\mathbb{N}}\subset L^{p}(\Omega)$ is such that
$\sup_{n\in\mathbb{N}}\left((1-s_{n})[\tilde{u}^{n}]^{p}_{s_{n},p}+\|u^{n}\|_{L^{p}(\Omega)}^{p}\right)<\infty,$ (3.34)
then
$\lim_{n\to\infty}(1-s_{n})\big|[\tilde{u}^{n}]^{p}_{s_{n},p,f_{n}}-[\tilde{u}^{n}]^{p}_{s_{n},p,f}\big|=0.$ (3.35)
Proof.
Since we can estimate
$\big|[\tilde{u}^{n}]^{p}_{s_{n},p,f_{n}}-[\tilde{u}^{n}]^{p}_{s_{n},p,f}\big|$ $\leq[\tilde{u}^{n}]^{p}_{s_{n},p}\sup_{(x,y)\in\mathbb{R}^{2d}}|f_{n}(x)f_{n}(y)-f(x)f(y)|$
[…]
the conclusion immediately follows from (3.34). ∎
Lemma 3.9.
If $(u^{n})_{n\in\mathbb{N}}\subset L^{p}(\Omega)$ is such that
$\sup_{n\in\mathbb{N}}\left((1-s_{n})[\tilde{u}^{n}]^{p}_{s_{n},p,f_{n}}+\|u^{n}\|_{L^{p}(\Omega)}^{p}\right)<\infty,$ (3.36)
then
$\sup_{n\in\mathbb{N}}(1-s_{n})[\tilde{u}^{n}]^{p}_{s_{n},p}+\|u^{n}\|_{L^{p}(\Omega)}^{p}<\infty.$
Proof.
On the one hand, by the inequality $(a+b)^{p}\leq 2^{p}(a^{p}+b^{p})$ for $a,b\geq 0$, exploiting symmetry, and passing to the $d$-dimensional spherical coordinates, we can estimate
$\begin{split}\int_{\mathbb{R}^{d}}\int_{B_{1}(y)^{c}}&\frac{|\tilde{u}^{n}(x)-\tilde{u}^{n}(y)|^{p}}{\|x-y\|^{d+s_{n}p}}\,\mathrm{d}x\,\mathrm{d}y\leq 2^{p}\int_{\mathbb{R}^{d}}\int_{B_{1}(y)^{c}}\frac{|\tilde{u}^{n}(x)|^{p}+|\tilde{u}^{n}(y)|^{p}}{\|x-y\|^{d+s_{n}p}}\,\mathrm{d}x\,\mathrm{d}y\\ &\leq 2^{p+1}\int_{\mathbb{R}^{d}}|\tilde{u}^{n}(y)|^{p}\int_{B_{1}(y)^{c}}\frac{\mathrm{d}x}{\|x-y\|^{d+s_{n}p}}\,\mathrm{d}y\\ &=2^{p+1}\|u^{n}\|^{p}_{p}\int_{1}^{\infty}\frac{d\omega_{d}}{\varrho^{1+s_{n}p}}\,\mathrm{d}\varrho=\frac{C(d,p)\|u^{n}\|^{p}_{p}}{s_{n}}\leq C(d,p)\|u^{n}\|^{p}_{p}\end{split}$ (3.37)
for all $n\in\mathbb{N}$, since $s_{n}\to 1^{-}$, thus we can assume $s_{n}>\frac{1}{2}$. On the other hand, recalling the notation in (2.1), since $\tilde{u}^{n}(y)$ is supported on $\Omega$, then $\tilde{u}^{n}(x)$ for $x\in B_{1}(y)$ is supported on $\Omega_{1}$, and since $f_{n}\to f$ uniformly and $f>0$, we have
$\begin{split}\int_{\mathbb{R}^{d}}\int_{B_{1}(y)}&\frac{|\tilde{u}^{n}(x)-\tilde{u}^{n}(y)|^{p}}{\|x-y\|^{d+s_{n}p}}\,\mathrm{d}x\,\mathrm{d}y=\int_{\Omega_{1}}\int_{\Omega_{1}}\frac{|\tilde{u}^{n}(x)-\tilde{u}^{n}(y)|^{p}}{\|x-y\|^{d+s_{n}p}}\,\mathrm{d}x\,\mathrm{d}y\\ &\leq\int_{\Omega_{1}}\int_{\Omega_{1}}\frac{|\tilde{u}^{n}(x)-\tilde{u}^{n}(y)|^{p}}{\|x-y\|^{d+s_{n}p}}\frac{f_{n}(x)f_{n}(y)}{\inf_{(x,n)\in\Omega_{1}\times\mathbb{N}}f_{n}^{2}(x)}\,\mathrm{d}x\,\mathrm{d}y\\ &=\frac{[\tilde{u}^{n}]^{p}_{s_{n},p,f_{n}}}{\inf_{(x,n)\in\Omega_{1}\times\mathbb{N}}f_{n}^{2}(x)}=C(\Omega_{1},f_{n})[\tilde{u}^{n}]^{p}_{s_{n},p,f_{n}}.\end{split}$ (3.38)
Since $f_{n}\to f$ uniformly and $f>0$, the constant $C$ can be made independent of $n$, and only dependent on the uniform limit $f$. The conclusion hence follows by combining (3.37) and (3.38), multiplying by $(1-s_{n})$, and owing to the assumption (3.36). We omit the plain details. ∎
4. Proof of Theorem˜1.2
4.1. Stability of Hilbertian gradient flows
We briefly recall some abstract machinery from [CdLKNP23] concerning Hilbertian gradient flows.
Let $\mathscr{H}$ be a Hilbert space endowed with scalar product $\langle\cdot,\cdot\rangle_{\mathscr{H}}$ and norm $\|\cdot\|_{\mathscr{H}}$. Given :($\mathscr{F}\colon\mathscr{H}\to$(-$\infty$$\infty$], we let $\mathscr{D}(\mathscr{F})=\{\,x\in\mathscr{H}\colon\mathscr{F}(x)<+\infty\,\}$ and
$\partial\mathscr{F}(x)=\left\{\,v\in\mathscr{H}\colon\liminf_{y\to x}\frac{\mathscr{F}(y)-\mathscr{F}(x)-\langle v,y-x\rangle_{\mathscr{H}}}{\|y-x\|_{\mathscr{H}}}\geq 0\,\right\}$
be the subdifferential of $\mathscr{F}$ at $x\in\mathscr{D}(\mathscr{F})$.
We recall the following result, which is a particular case of [CdLKNP23]*Prop. 3.7. Here and below, given any vector space $\mathscr{V}$, we let $\mathscr{V}^{*}$ be the algebraic dual space of $\mathscr{V}$ and $\mathscr{V}^{\prime}$ the topological dual space of $\mathscr{V}$.
Proposition 4.1.
Let $\mathscr{K}$ be a dense subspace of $\mathscr{H}$. Let :($\mathscr{F}\colon\mathscr{H}\to$(-$\infty$$\infty$] be a proper, convex, and strongly lower semicontinuous functional and let $x\in\mathscr{D}(\mathscr{F})$. If there exists $T\in\mathscr{K}^{*}$ such that
$\lim_{t\to 0}\frac{\mathscr{F}(x+ty)-\mathscr{F}(x)}{t}=T(y),\quad\textrm{for every }y\in\mathscr{K}\,$,
then either $\partial\mathscr{F}(x)=\emptyset$ or $\partial\mathscr{F}(x)=\{v\}$, where $v$ is the (unique) element in $\mathscr{H}$ satisfying $T(y)=\langle v,y\rangle_{\mathscr{H}}$ for every $y\in\mathscr{K}$. In particular, $T\in\mathscr{K}^{\prime}$ and $v$ is its unique continuous extension to $\mathscr{H}^{\prime}$.
We also recall the following stability result, which is contained in [CdLKNP23]*Thm. 3.8.
Theorem 4.2.
[…]
  1. (a)
    $(\mathscr{F}_{n})_{n\in\mathbb{N}}$ $\Gamma$-converges to some proper functional :($\mathscr{F}_{\infty}\colon\mathscr{H}\to$(-$\infty$$\infty$] with respect to the strong $\mathscr{H}$-convergence;
  2. (b)
    any bounded sequence $(x_{n})_{n\in\mathbb{N}}\subset\mathscr{H}$ such that $\sup_{n\in\mathbb{N}}\mathscr{F}_{n}(x_{n})<\infty$ admits a strongly $\mathscr{H}$-convergent subsequence.
If $(x^{n}_{0})_{n\in\mathbb{N}}\subset\mathscr{H}$ is such that $x^{n}_{0}\in\mathscr{D}(\mathscr{F}_{n})$ for every $n\in\mathbb{N}$, $\sup_{n\in\mathbb{N}}\mathscr{F}_{n}(x^{n}_{0})<\infty$ and $x^{n}_{0}\to x^{\infty}_{0}$ strongly in $\mathscr{H}$ for some $x^{\infty}_{0}\in\mathscr{H}$, then the following hold:
  1. (i)
    $x^{\infty}_{0}\in\mathscr{D}(\mathscr{F}_{\infty})$;
  2. (ii)
    for every $T>0$, the problem
    $\begin{cases}\dot{x}(t)\in-\partial\mathscr{F}_{n}(x(t)),\quad\text{for a.e.\ }t\in(0,T)\,,\\[4.30554pt] x(0)=x^{n}_{0},\end{cases}$
    admits a unique solution $x_{n}\in H^{1}([0,T];\mathscr{H})$ for every $n\in\mathbb{N}\cup\{\infty\}$;
  3. (iii)
    $(x_{n})_{n\in\mathbb{N}}$ weakly converges to $x_{\infty}$ in $H^{1}([0,T];\mathscr{H})$.
Moreover, if $\lim_{n\to\infty}\mathscr{F}^{n}(x^{n}_{0})=\mathscr{F}_{\infty}(x^{\infty}_{0})$, then actually $(x_{n})_{n\in\mathbb{N}}$ strongly converges to $x_{\infty}$ in $H^{1}([0,T];\mathscr{H})$ and also
$x_{n}(t)\xrightarrow{\mathscr{H}}x_{\infty}(t)\quad\text{and}\quad\mathscr{F}_{n}(x_{n}(t))\to\mathscr{F}_{\infty}(x_{\infty}(t))\quad\text{for every}\ t\in[0,T].$
4.2. Proof of Theorem˜1.2
The validity of Theorem˜1.2 follows by combining the abstract results above with the following proposition. Here and below, $\langle\cdot,\cdot\rangle$ denotes the standard scalar product in $L^{2}(\mathbb{R}^{d})$.
Proposition 4.3.
Let $\varphi\in C_{c}^{\infty}(\Omega)$ and $u\in L^{2}(\Omega)$. The following hold:
  1. (i)
    if $u\in H^{1}_{0}(\Omega)$, then
    $\lim_{t\to 0}\frac{\|\nabla(u+t\varphi)\|^{2}_{2,f^{2}}-\|\nabla u\|_{2,f^{2}}^{2}}{t}=\langle(-\mathfrak{D})^{f}u,\varphi\rangle;$ (4.1)
  2. (ii)
    if $\tilde{u}]_{s,2,f}<\infty$ for some $s\in(0,1)$, then
    $\lim_{t\to 0}\frac{[u+t\varphi]_{s,2,f}^{2}-[u]_{s,2,f}^{2}}{t}=\langle(-\mathfrak{D})^{s,f}u,\varphi\rangle.$ (4.2)
Proof.
We only prove (4.2), the proof of (4.1) being straightforward. We note that
$[u+t\varphi]_{s,2,f}^{2}$ $=[u]_{s,2,f}^{2}+t^{2}[\varphi]_{s,2,f}^{2}$
$\quad+2t\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{(u(x)-u(y))(\varphi(x)-\varphi(y))}{\|x-y\|^{d+2s}}f(x)f(y)\,\mathrm{d}x\,\mathrm{d}y$
for every $t\in\mathbb{R}$, and thus we easily get that
$\lim_{t\to 0}\frac{[u+t\varphi]_{s,2,f}^{2}-[u]_{s,2,f}^{2}}{t}$ $=2\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{(u(x)-u(y))(\varphi(x)-\varphi(y))}{\|x-y\|^{d+2s}}f(x)f(y)\,\mathrm{d}x\,\mathrm{d}y$
$=4\int_{\mathbb{R}^{d}}u(x)f(x)\lim_{r\to 0^{+}}\int_{\mathbb{R}^{d}\setminus B_{r}}\frac{\varphi(x)-\varphi(y)}{\|x-y\|^{d+2s}}f(y)\,\mathrm{d}y\,\mathrm{d}x$
$=\langle u,(-\mathfrak{D})^{s,f}\varphi\rangle=\langle(-\mathfrak{D})^{s,f}u,\varphi\rangle$
in virtue of the distributional definition in (1.7), concluding the proof. ∎
We are now ready to prove Theorem˜1.2.
Proof of Theorem˜1.2.
Since $f_{n}\to f$ uniformly and $f>0$, the functionals given by $\mathscr{F}_{n}(u)=(1-s_{n})[\tilde{u}]_{s_{n},2,f_{n}}^{2}$ are positive for $n\gg 1$. Further, they are easily shown to be convex. By Theorem˜1.1(ii)(iii), $(\mathscr{F}_{n})_{n\in\mathbb{N}}$ $\Gamma$-converges to the functional $\mathscr{F}(u)=K_{d,2}\|\nabla u\|^{2}_{2,f^{2}}$ in the strong topology of $L^{2}(\Omega)$, whereas by Theorem˜1.1(i) every sequence $(u_{n})_{n\in\mathbb{N}}\subset L^{2}(\Omega)$ such that $\sup_{n\in\mathbb{N}}\mathscr{F}_{n}(u_{n})<\infty$ admits a strong $L^{2}$-limit $u\in L^{2}(\Omega)$. The conclusion hence follows by Theorems˜4.2 and 4.1 combined with Proposition˜4.3. ∎